40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Mar 22 16:29:04 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
    Reply  Thu Mar 24 19:07:38 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
       Reply  Sat Mar 26 13:36:09 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT IMG_2468.JPGIMG_2466.JPG
          Reply  Tue Mar 29 23:23:37 2016, rana, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
             Reply  Wed Mar 30 08:29:15 2016, tara, DailyProgress, PMC, test on new PMC IMG_2528.JPGIMG_2540.JPG
                Reply  Wed Mar 30 10:28:44 2016, Evan, DailyProgress, PMC, test on new PMC 
                   Reply  Thu Mar 31 15:11:27 2016, tara, DailyProgress, PMC, test on new PMC 
                      Reply  Tue Apr 5 16:32:36 2016, tara, DailyProgress, PMC, test on new PMC 
                         Reply  Wed Apr 6 15:02:50 2016, tara, DailyProgress, PMC, test on new PMC 
                            Reply  Fri Apr 8 09:44:49 2016, tara, DailyProgress, PMC, test on new PMC IMG_2634.JPGIMG_2635.JPGIMG_2633.JPG
Message ID: 1623     Entry time: Tue Apr 5 16:32:36 2016     In reply to: 1620     Reply to this: 1624
Author: tara 
Type: DailyProgress 
Category: PMC 
Subject: test on new PMC 

We are hacking the PMC card, to try to lock the PMC without using a computer.

Aidan helped me try to lock the PMC with the current equipment (LIGO PMC card, the schematic can be found here). Since, the computer for controlling the card is not here and we cannot put it back together within one week I have left, we are trying to use the card and send in signals to control the card manually. For today test, we used a spare PMC left on the crate (I think Frank used it to drive the PZT for shaking the table, long time ago. It might be modifed from what we saw on the schematic.)

We tested the control signal for the high voltage output using the following procedures.

  • Power up the card with +/- 24 V with a power supply
  • Use Acromag to send Vin ( I think it can send in +/- 10 V, but we used only [-6 , 2] V) .
  • Use HV from the power supply on the rag with 0-200 V. 
  • Measure Vout at J6 (where it will be connected to the PMC's PZT)

The HV drive is working and linear with the input control [HV = -24.1*Vin + 40] (Vin = [-6,2] V, HV = [4,185] V). At first, we injected the signal at INOFFSET2 (through R4), but we couldn't see the signal out. I'm not sure why, but we are working on it.

To do:

  • Setup RFPD on the table for the PMC
  • Add the EOM on the path (may be temporary, as the mount is not ready)
  • Connect the LO to EOM and the PMC card
  • Check Error signal
  • Lock the cavity
  • Note: the mysterious mode I reported in previous log may be mode hopping (as suggested by Johanness during Pizza meeting) I'll check that by operate the laser at different slow offset and see if there's any change or not
ELOG V3.1.3-